OpenStudy (anonymous):

express y = 3/4x + 6 in standard form.

OpenStudy (anonymous):

Standard for is Ax + By + C = 0 So just bring everything to one side so it equals 0

OpenStudy (anonymous):

i thought standard form was Ax + By = C,,,

OpenStudy (anonymous):

It's not.

OpenStudy (anonymous):

ok so how do i move the fraction? subtraction like usual?

OpenStudy (anonymous):

Yes, that's it.

OpenStudy (anonymous):

so the answer is 3/4x +y-6=0?

OpenStudy (anonymous):

3/4x-y +6 = 0 OR -3/4x +y-6=0

OpenStudy (anonymous):

that doesn't seem right,,,

OpenStudy (anonymous):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.

OpenStudy (anonymous):

Ax + By + C = 0 still what I see. http://www.mathsisfun.com/algebra/standard-form.html

OpenStudy (anonymous):

sigh,,, that is general form from what that page said,,, ty for your help ???

OpenStudy (anonymous):

Can u draw the question

OpenStudy (anonymous):

y=3/4x+6 Ax +By=C

OpenStudy (anonymous):

I guess its Ax+By +C=0 is the standard form for the quadratic equation

OpenStudy (anonymous):

no that's not it

OpenStudy (anonymous):

Ax + By = C is standard form. Ax + By + C = 0 is sometimes called standard form but is actually general form.